The first painter can paint a room in 6 hours, and the second can paint the same room in 8 hours.How long will it take if they work together?
Answer by jorel1380(3719)

1/6+1/8=1/x
4x+3x=24
7x=24
x=24/7
It takes 24/7 hours working together to paint the room..